# TWIST: Training-free and Label-free Short Text Clustering through Iterative Vector Updating with LLMs

### v1

@@ Abstract
In this paper, we propose a training-free and label-free method for short text clustering that can be used on top of any existing embedder. In the context of customer-facing chatbots, companies are dealing with large amounts of user utterances that need to be clustered according to their intent. In these commercial settings, no labeled data is typically available, and the number of clusters is not known. Our method is based on iterative vector updating: it constructs sparse vectors based on representative texts, and then iteratively refines them through LLM guidance. Our method achieves comparable or superior results to state-of-the-art methods that use contrastive learning, but without assuming prior knowledge of clusters or labels. Experiments on diverse datasets and smaller LLMs show that our method is model agnostic and can be applied to any embedder, with relatively small LLMs, and different clustering methods. We also show that our method scales to large datasets, reducing the computational cost of the LLM. These low-resource, adaptable settings and the scalability of our method make it more aligned with real-world scenarios than existing clustering methods. Our source code is available here: https://anonymous.4open.science/r/sparse_vector-F4C9/README.md
@@ 1 Introduction
The clustering task is often approached as a Generalized Category Discovery (GCD) scenario (Vaze et al., 2022), where limited labeled data and substantial unlabeled data are used to cluster all unlabeled samples, which include both known and unknown classes. Recently, other studies have explored an emerging data scenario, where no labels are provided at all (De Raedt et al., 2023; Viswanathan et al., 2024b; Zhang et al., 2023; Lin et al., 2025). This scenario reduces reliance on human annotation and shows greater flexibility.
A common approach for the emerging data scenarios is to train an embedder with contrastive learning to improve text representations (Zhang et al., 2021). With the advancement of LLMs, training-free approaches have also been proposed (De Raedt et al., 2023). The emerging data scenario has attracted increasing attention in recent research (Zhang et al., 2023; Feng et al., 2024; Lin et al., 2025). Despite the progress of prior work, these label-free approaches still face certain limitations and leave room for further improvement. First, methods involving LLM feedback often assume known number of clusters during method development (Zhang et al., 2023; Feng et al., 2024), which is difficult to estimate in real-world settings. Second, training-free methods have mainly been explored with large, proprietary models (Viswanathan et al., 2024b; De Raedt et al., 2023), which can be costly. Third, the scalability of these methods remains a concern. Motivated by these challenges, we propose a training-free approach that requires only a smaller LLM, alleviates the known-cluster assumption, and is scalable. The core idea of our method is to construct new representations (referred to as sparse vectors in this paper) such that texts belonging to the same cluster become more similar. Our method consists of two stages: we first construct sparse representations based on representative texts, and then iteratively refine them by combining LLM guidance with pre-trained embeddings.
The contributions of our method are as follows:
  • Inclusive and easy to optimize: Builds on existing embedders and requires only some hyperparameter tuning.
  • Flexible: Does not assume a known number of clusters, nor require labeled data, contrastive learning on the task specific datasets, or fine-tuning, enabling easier adaptation to new domains.
  • Low-resource: Even with small LLMs and embedders, the results are already competitive with methods that rely on labeled data or fine-tuning.
  • Scalable: Does not require LLM feedback for every text, unlike many previous studies.
@@ 2 Related Work
Generalized Category Discovery (GCD) is often studied for this task (Liang et al., 2024b; Liang and Liao, 2023; Zou et al., 2025b). These works use limited labeled text and substantial unlabeled text to learn clustering representations. However, these studies use train–test splits within a single dataset, which implicitly assumes that new categories are similar to the known ones. In practice, emerging data may show topic drift. Recently, with the advancement of LLMs, many studies have explored the emerging data scenario (Zhang et al., 2023; Viswanathan et al., 2024b; Lin et al., 2025), where no labels are available. We follow this unsupervised scenario.
Boosting Unsupervised Short Text Clustering with LLM-Guided Contrastive Learning
Contrastive learning is commonly used in unsupervised short text clustering (Zhang et al., 2021, 2023; Liang et al., 2024a). Zhang et al. (2023) use LLMs to construct hard triplets to derive training pairs, which are then used to fine-tune a small embedder. Liang et al. (2024a) use LLMs to refine texts by generating descriptors, then using small language models for contrastive learning to align the embeddings of texts and their descriptors. However, these methods either rely on knowing the number of clusters, proprietary models, or fine-tuning. Our method does not have these requirements.
Training-Free Methods for Unsupervised Short Text Clustering
Training-free approaches use a frozen pre-trained embedder: De Raedt et al. (2023) select prototypical utterances and then use an LLM to generate a set of descriptive utterance labels for these prototypes. The labels and texts are then encoded to form better representations. Viswanathan et al. (2024a) use LLMs to improve texts by generating key phrases for each text, which are then separately encoded into embeddings and concatenated for clustering. Lin et al. (2025) use LLMs for selection rather than text enrichment. They first encode all of the texts, and then use the LLM to select similar pairs for each text. Each text embedding is then obtained by average pooling the embeddings of the text and its similar pairs. These methods have in common that they refine original embeddings to improve clustering. In contrast, we construct a new vector using LLM feedback. Feng et al. (2024) use LLMs to refine edge points. However, the assumption of a known number of clusters can still be a core problem in real-world applications.
3 Theoretic formulation and assumptions
@@ 3.1 Problem definition
In our work, we consider two variations of the short text clustering task: (i) the emerging data scenario and (ii) the GCD scenario. In the emerging data scenario (Lin et al., 2025), the only available resource is a collection of unlabeled texts $D^{test}\triangleq\{x_{i}\}_{i=1}^{N}$, where each $x_{i}\in D^{test}$ corresponds to a short text and $N$ is the total number of samples. The dataset is used as a whole for clustering and evaluation, without any train-test split. In contrast, the GCD scenario has access to additional data, typically including some labeled samples (Zou et al., 2025b). In this work, we adapt this setting by considering only an additional unlabeled dataset $D^{u}\triangleq\{x_{i}\}_{i=1}^{N_{u}}$, in line with the goal of low-resource. The GCD scenario is included for the purpose of testing whether extra data improves clustering and generalizability of our method.
For both scenarios, the task is to partition $D^{test}$ into clusters $\{\hat{S}_{l}\}_{l=1}^{\hat{K}}$, where $\hat{K}$ is the number of predicted clusters. Note that $\sum_{l=1}^{\hat{K}}|\hat{S}_{l}|=N$. The goal is to ensure that the clustering results $\{\hat{S}_{l}\}_{l=1}^{\hat{K}}$ are as close as possible to the true partition $\{S_{l}\}_{l=1}^{K}$, where $K$ is the number of unique clusters in the ground truth. Importantly, unlike many existing studies (De Raedt et al., 2023; Feng et al., 2024; Zou et al., 2025b), which assume that $K$ is known in advance, we make no such assumption.
@@
3.2 Constructing the Sparse Vector
Basic Assumption
Considering the fact that texts belonging to the same cluster are similar to each other, we represent each text as a sparse bag-of-texts vector that directly encodes its similarity to other texts. Formally, for any $x_{i},x_{j}\in D$ with $i\neq j$:
  • $x_{i}$ and $x_{j}$ are similar if $x_{i},x_{j}\in S_{k}$ for some cluster $S_{k}$.
  • $x_{i}$ and $x_{j}$ are not similar if $x_{i}\in S_{k}$ and $x_{j}\in S_{l}$ with $k\neq l$.
Construction of Bag-of-Texts
Since similarity is determined by clusters, it is not necessary to use all texts in $D$ to represent sparse vectors; instead, a representative subset of texts can constitute our bag of texts, denoted as $D^{r}\subseteq D$, where $|D^{r}|\triangleq d$. The set of representative texts $D^{r}$ should contain at least one element from each cluster. That is, for any cluster $S\in\{S_{l}\}^{K}_{l=1}$, there exists at least one text $x\in D^{r}$ such that $x\in S$.
Therefore, to cover all possible clusters, the minimum size of $d$ is K1K-1. Since $K$ is unknown, the most practical approach is to choose a relatively large $d$. For each $x\in D$, we define a $d$-diemnsional vector $\mathbf{z}$ assigned to the representative texts, where the $j$-th coordinate takes the value 1 if $x$ is similar to the $j$-th representative text, and 0 otherwise.
@@ 3.2 Constructing the Sparse Vector
Given the assumption, for any $x_{i},x_{j}\in D$ with $i\neq j$ that belong to the same cluster $S_{k}$, the corresponding sparse vectors $\mathbf{z}_{i}$ and $\mathbf{z}_{j}$ will be identical.
The proposition is straightforward, so we omit the proof.
Adaptation to real-world practice
Although the proposition holds in the idealized setting, in practice similarity is typically determined using LLMs, which may introduce errors, and scanning the entire dataset for each $x$ is computationally expensive. To address these, we construct a sparse vector $\mathbf{z}$ in a way that is adaptable to real-world practice: for each $x\in D$, instead of assigning fixed value in $\{0,1\}$ based on representative texts, we propose an updating procedure that accounts for uncertainty. In addition, to address limited computational resources, we use an existing embedder to identify a relatively small set of $m$ similar candidates that are likely to belong to the same category, as empirically shown in previous research (Lin et al., 2025; Zou et al., 2025a). The details of the method will be discussed in the next section.
@@ 4 Computational Method
Our method aims to construct sparse vectors that capture the similarity of texts within the same cluster. It consists of two stages. In the initial stage, we select representative texts, which are then used to construct a sparse vector for each text. In the second stage, we iteratively update each text’s sparse vector by considering its similar texts until convergence, illustrated in Figure 1.
Unlike many prior studies that assume the number of clusters $K$ is known, we consider a more practical setting in which $K$ is unknown, and thus approximate the true number of clusters by the number of representative texts, $d$. This is less challenging than estimating $K$, as it only requires ensuring that $d$ is large enough to cover $K$ clusters.
Figure 1: Construction of Sparse vectors. Initial stage: We first encode each text in $D$ using a pretrained embedder and partition them into $d$ clusters. We then select the medoid of each cluster as the representative text and construct one-hot vectors for these representatives. The remaining text vectors are obtained based on LLM selection. Iterative stage: We update all of the vectors iteratively until convergence.
4.1 Initial stage: construction of sparse representation
To build sparse vector representations, we first need to create a set of representative samples $D^{r}$. To this end, we first encode each text using an existing embedder $F$ to obtain pre-trained embeddings:
$\mathbf{x}_{i}=F(x_{i}).$
Next, we apply a clustering algorithm to partition these embeddings into $d$ clusters.
We use agglomerative clustering (Murtagh and Legendre, 2014) as it produces deterministic results, which facilitates reproducibility. For each initial cluster , we derive the medoid. We refer to the set of medoids as the representative text, denoted as:
$D^{r}\triangleq\{x_{c_{1}},\dots,x_{c_{d}}\}.$
and let $\mathbf{X}^{r}$ be $D^{r}$ corresponding pre-trained embeddings, where $\mathbf{X}^{r}=\{\mathbf{x}_{i}:x_{i}\in D^{r}\}$. We then transform the representative texts into one-hot vectors, which serve as the basis vectors for the sparse representation.
For the remaining texts $x_{j}\in D\setminus D^{r}$, we construct their sparse vectors using this basis. For each $x_{j}$, we compute the pairwise cosine distances between its embedding $\mathbf{x}_{j}$ and the medoid embeddings in $\mathbf{X}^{r}$. By sorting these distances, we select the top $m$ closest representative texts to $x_{j}$, denoted as $x_{j1},\dots,x_{jm}$. Then, we use an LLM to identify the similar text(s) to $x_{j}$ among these representative texts, denoted as $M_{j}\triangleq LLM(x,x_{j1},\dots,x_{jm})\in\mathcal{P}(\{x_{j1},\dots,x_{jm}\})\cup\{\text{None}\}$, where $\mathcal{P}(\cdot)$ represents the power set, i.e., all possible subsets of the candidate texts, and ‘None’ indicates that no representative text is selected. If the LLM output is not None, the sparse vector of $x_{j}$ is then derived by averaging the sparse vectors of the texts in $M_{j}$, denoted as $\mathbf{z}_{j}$. Instead of assigning a fixed value of 1 to the selected dimension, we average the vectors to capture how strongly each representative text contributes to the feature. We then apply L2 normalization to ensure a consistent scale for the iterative stage. If the LLM output is None, we make this text into a representative text itself and increase the dimensionality of all sparse vectors by one and construct its corresponding one-hot vectors.
Note: if we set $d=N$, all texts will be representative texts and one-hot encoded.
4.2 Iterative stage: sparse vectors refinements
In this stage, we update the sparse vectors of all texts using the dynamic subset of $m$ candidates.
At each iteration step $t$ (with the initial stage being $t=0$), for each $x_{j}\in D$, we first concatenate its pre-trained embedding and sparse vector: $\mathbf{u}_{j}^{t}\triangleq\mathbf{x}_{j}\oplus\mathbf{z}_{j}^{t}$. We then compute the pairwise cosine distances between $\mathbf{u}_{j}^{t}$ and the concatenated representations of the all texts $\mathbf{U}^{t}$. By sorting these distances, we select the top $m$ closest texts to $x_{j}$, denoted as $x^{t}_{j1},\dots,x^{t}_{jm}$. The concatenation ensures that the pre-trained embedding guides the sparse vectors. Also, the reason we do iterative updating instead of just once is to gain more stable sparse vectors.
Next, we use an LLM to identify the subset of the texts most similar to $x_{j}$, denoted as $M^{t}_{j}\triangleq LLM(x_{j},x^{t}_{j1},\dots,x^{t}_{jm})\in\mathcal{P}(\{x^{t}_{j1},\dots,x^{t}_{jm}\})\cup\{\text{None}\}$. If the LLM output is None, we fix sparse vectors, i.e., $\mathbf{z}_{j}^{t+1}=\mathbf{z}_{j}^{t}$. Otherwise, we update the sparse vectors by averaging the sparse vector of the texts in $M^{t}_{j}$. We then apply L2 normalization, as explained previously.
For $x_{j}$, the update process is ended if the sparse vector at iteration $t+1$ is almost identical to those at iteration $t$, as we define by a cosine similarity greater than 0.99. For efficiency, we set $m=30$ and cap the maximum number of iterations at 10.

### v2

# LLMs Enable Bag-of-Texts Representations for Short-Text Clustering

@@ Abstract
In this paper, we propose a training-free method for unsupervised short text clustering that relies less on careful selection of embedders than other methods. In customer-facing chatbots, companies are dealing with large amounts of user utterances that need to be clustered according to their intent. In these settings, no labeled data is typically available, and the number of clusters is not known. Recent approaches to short-text clustering in label-free settings incorporate LLM output to refine existing embeddings. While LLMs can identify similar texts effectively, the resulting similarities may not be directly represented by distances in the dense vector space, as they depend on the original embedding. We therefore propose a method for transforming LLM judgments directly into a bag-of-texts representation in which texts are initialized to be equidistant, without assuming any prior distance relationships. Our method achieves comparable or superior results to state-of-the-art methods, but without embeddings optimization or assuming prior knowledge of clusters or labels. Experiments on diverse datasets and smaller LLMs show that our method is model agnostic and can be applied to any embedder, with relatively small LLMs, and different clustering methods. We also show how our method scales to large datasets, reducing the computational cost of the LLM use. The flexibility and scalability of our method make it more aligned with real-world training-free scenarios than existing clustering methods. Our source code is available here: https://anonymous.4open.science/r/BoT_vector-2E0C/README.md
@@ 1 Introduction
Figure 1: illustration of our proposed method. Left: traditional approach using the LLM output to refine the embeddings space. Right: our approach using the LLM output to directly construct a bag-of-texts space.
@@
The short-text clustering task is often approached as a Generalized Category Discovery (GCD) scenario (Vaze et al., 2022), where limited labeled data and substantial unlabeled data are used to cluster all unlabeled samples, which include both known and unknown classes. Recently, other studies have explored an emerging data scenario, where no labels are provided at all (De Raedt et al., 2023; Viswanathan et al., 2024b; Zhang et al., 2023; Lin et al., 2025). This scenario reduces reliance on human annotation and shows greater flexibility.
Regardless the setting, the classic approach is to train an embedder with contrastive learning (Zhang et al., 2021). In recent years, most methods have been designed to use LLMs to guide the optimization (Zhang et al., 2023; Zou et al., 2025b). Many LLM-guided training-free approaches have been proposed as well to enable easier domain adaptation and reduce the need for fine-tuning (De Raedt et al., 2023; Viswanathan et al., 2024b; Lin et al., 2025). Although these training-free approaches lead to substantial improvement, they only partially capture LLM-derived similarities and remain heavily dependent on the original embeddings. The resulting representation may still not reflect LLM preference, despite this being a strong signal. Hence our objective: to construct text representations that directly incorporate LLM similarity estimations into the vector space.
The core idea of our method (Figure 1) is to construct new representations (referred to as bag-of-texts vectors) such that texts belonging to the same cluster get more similar representations. Our method consists of two stages: we first construct bag-of-texts (BoT) representations based on representative texts, and then iteratively refine them by LLM guidance. In summary, the contributions of our method are as follows:
  • A Strategy for Encoding LLM Preferences Unlike prior work that refines existing embeddings, we introduce a Bag-of-Texts (BoT) representation for encoding LLM outputs.
  • Flexible: Our method can be added to any embedder and does not require fine-tuning, enabling easier adaptation to new domains.
  • Low-resource: Our method reduces the need for human labeling, can be implemented using relatively small, zero-shot models, and achieves results competitive with labeled or fine-tuned approaches.
@@ 2 Related Work
Generalized Category Discovery (GCD) is often studied for this task (Liang et al., 2024b; Liang and Liao, 2023; Zou et al., 2025b). These works use limited labeled text and substantial unlabeled text to learn clustering representations. However, these studies use train–test splits within a single dataset, which implicitly assumes that new categories are similar to the known ones. In practice, emerging data may show topic drift. Recently, with the advancement of LLMs, many studies have explored the emerging data scenario (Zhang et al., 2023; Lin et al., 2025), where no labels are available. We follow this unsupervised scenario to support real world applications.
LLM-Guided Contrastive Learning for Short Text Clustering
With the advancement of LLM, recent work has begun to incorporate LLM feedback into the optimization process. Zhang et al. (2023) use LLMs to construct hard triplets to derive training pairs. Liang et al. (2024b) identify samples in ambiguous regions and use LLMs to reassign them to new clusters, which are then used as training data. Zou et al. (2025b) proposed a unified framework that actively learns from diverse and quality-enhanced LLM feedback, including instance-level, category description, and uncertain-instance alignment. However, these methods either rely on knowing the number of clusters, proprietary models, or fine-tuning. Our method does not have these requirements.
LLM-Guided Training-Free Approaches for Short Text Clustering
Training-free approaches use a frozen pre-trained embedder and use LLMs in two ways. The first category is input text enrichment: De Raedt et al. (2023) select prototypical utterances and then use an LLM to generate a set of descriptive utterance labels for these prototypes. Viswanathan et al. (2024a) use LLMs for few-shot in-context learning to improve texts by generating key phrases for each text. The second category is embedding post-processing: Lin et al. (2025) use LLMs for selection rather than text enrichment. They first encode all of the texts, and then use the LLM to select similar pairs for each text. Each text embedding is then obtained by average pooling its similar pairs. These methods have in common that they refine original embeddings. In contrast, we construct a new vector using LLM feedback.
3 Bag-of-Texts Theoretical Justification
@@ 3.1 Problem Definition
The task is to partition $D^{test}$ into $K$ clusters. $K$ is treated as prior knowledge in most existing work (De Raedt et al., 2023; Zhang et al., 2023; Feng et al., 2024; Zou et al., 2025b), and many approaches incorporate $K$ when refining embeddings. Our proposed method does not make any assumption on the knowledge of $K$ thus can be implemented without knowing the number of clusters. For evaluation, however, we incorporate $K$ for fair comparison to previous methods.
In line with prior work, we address two scenarios: (i) the emerging data scenario and (ii) the Generalized Category Discovery (GCD) scenario. In the emerging data scenario (Zhang et al., 2023; Lin et al., 2025), the only available resource is a collection of unlabeled texts $D^{test}\triangleq\{x_{i}\}_{i=1}^{N}$, where each $x_{i}\in D^{test}$ corresponds to a short text and $N$ is the size of the test dataset. The dataset is used as a whole for clustering and evaluation, without any train-test split. In contrast, the GCD scenario typically has access to additional data, typically including some labeled examples (Liang et al., 2024b; Zou et al., 2025b). As opposed to prior work, to align the GCD scenario with the objective of low-resource settings, we restrict the availability of additional data to only an unlabeled dataset $D^{u}\triangleq\{x_{i}\}_{i=1}^{N_{u}}$.
@@
3.2 Embedding Refinement via Bag-of-Texts
Existing LLM-guided short-text clustering approaches often refine embeddings either directly using contrastive learning Zhang et al. (2023), indirectly by text enrichment De Raedt et al. (2023); Viswanathan et al. (2024b), or through embedding post-processing Lin et al. (2025). While contrastive optimization enforces inequality through gradient updates, it requires training on new samples, which makes it impractical for low-resource settings. Text enrichment approaches, on the other hand, can perform for low-resource scenarios but lack explicit guarantees for pulling similar texts closer and pushing dissimilar ones apart. Embedding post-processing approaches heavily depend on the embedder used.
Our novel bag-of-texts representation combines the strengths of these approaches and directly enforces inequality through text-based LLM outputs, without requiring training data or expensive gradient-based updates of embedders. Our approach mimics contrastive learning by reconstructing the vector space using a bag-of-texts representation and updating this representation using LLM guidance. Using bag-of-texts representation, all texts start with equivalent distances in the vector space, and LLM output about similarity of texts to each other is directly translated to the vector representation, moving each text toward the correct region of the space; see Figure 1 for illustration.
Formally, for a text $x$, let $l_{x,x^{+}}$ and $l_{x,x^{-}}$ denote similarities between $x$ and a positive and negative example, and $l^{{}^{\prime}}_{x,x^{+}}$ and $l^{{}^{\prime}}_{x,x^{-}}$ be the corresponding similarities after embedding refinement. The positive and negative examples are identified by a function (e.g., an LLM or a human), which determines whether the text $x$ belongs to the same cluster as example $x^{+}$ (positive) or to a different cluster from the example $x^{-}$ (negative). Ideally, our refinement approach should satisfy $l^{{}^{\prime}}_{x,x^{+}}>l^{{}^{\prime}}_{x,x^{-}}$. Below, we formally explain how our bag-of-texts representation achieves this goal.
Assumption 1.
We assume that a function (e.g., an LLM or a human) makes a selection based on the following rules.
  • Similarity: For any texts $x_{i}$ and $x_{j}$ with $i\neq j$, $x_{i}$ and $x_{j}$ are similar if they belong to the same cluster $S_{k}$, and not similar if they belong to different clusters $S_{k}$ and $S_{h}$ with $k\neq h$.
  • Distinct Clusters: Each text is assigned to exactly one cluster $S\in\mathcal{S}$.
Note that $\mathcal{S}$ in Assumption 1 refers to the clusters defined by the function; there is no knowledge of a ground truth. We assume distinct clusters as is conventional in prior work (An et al., 2024; Lin et al., 2025).
Definition 1 (Bag-of-Texts).
Let $g(x,D)$ be a function that takes a text $x$ and a set of texts $D$, and outputs a bag-of-text representation $\mathbf{z}\in\mathbb{R}^{|D|}$. In this representation, the entry corresponding to each text in $D$ that belongs to the same cluster as $x$ is assigned a value of 1, while all other entries are 0.
@@ 3.2 Embedding Refinement via Bag-of-Texts
Given Assumption 1, for any $x_{i},x_{j}\in D$ and their corresponding bag-of-texts representation $\mathbf{z}_{i},\mathbf{z}_{j}$, the function $g(x,D)$ guarantees that $l^{{}^{\prime}}_{x,x^{+}}>l^{{}^{\prime}}_{x,x^{-}}$.
The proof of the proposition builds on Definition 1 that if two texts $x_{i},x_{j}\in D$ belong to the same cluster, their bag-of-texts representations $\mathbf{z}_{i}$ and $\mathbf{z}_{j}$ are identical. Considering cosine similarity as an example similarity function, $\text{cosine}(\mathbf{z}_{i},\mathbf{z}_{j})=1$ if they are identical, and $\text{cosine}(\mathbf{z}_{i},\mathbf{z}_{j})=0$ if they are dissimilar.
3.3 Adaptation to Real-world Practice
Although the proposition holds in the idealized setting, in practice the similarity between texts is typically determined using LLMs, which may introduce errors, and scanning the entire dataset for each $x$ is computationally expensive. We outline the modifications needed for practical implementation. The detailed analysis is in appendix A.
Memory constraints.
Large datasets lead to high-dimensional bag-of-text vectors. To address these issues, we construct bag-of-texts vectors using a representative subset $D^{r}\subseteq D$ with $|D^{r}|\triangleq d$ , rather than all texts. This dimensionality reduction is possible because it relies on the assumption that texts within the same cluster are similar; therefore, it is not necessary to use all texts in $D$ to represent bag-of-texts vectors. Instead, the bag-of-texts vectors of the remaining texts can be represented via the vectors of the selected representatives.
Computation constraints.
Selecting similar texts requires scanning the entire dataset for each text, which is computationally expensive with an LLM. To reduce computation, we use a pre-trained embedder to retrieve a small set of $m$ candidate texts that are likely to be similar, as supported by prior work (Lin et al., 2025; Zou et al., 2025a).
Selection uncertainty.
During automatic selection (e.g., by LLM) inconsistent preferences can occur and violate Assumption 1. For example, $(x_{i},x_{j})$ may be identified as belonging to the same cluster, $(x_{i},x_{l})$ as different, yet $(x_{j},x_{l})$ as the same. Also, in practice, each selection is made with some inherent confidence. To reduce these inconsistencies, we perform iterative update and assign the mean of the selected bag-of-texts vectors rather than a fixed value of 1.
@@ 4 Computational Method
Our method aims to construct Bag-of-Texts (BoT) vectors that capture the similarity of texts within the same cluster. It consists of two stages. First, we select representative texts, which are then used to construct a BoT vector for each text. Second, we iteratively update each text’s BoT vector by considering its similar texts until convergence.
4.1 Initial stage: construction of BoT representation
Algorithm 1 Select Representative Texts
0:$D$, $\mathbf{X}$, representative texts size $d$
[…]
3:for $j=1,\dots,d$ do
4:  $\mathbf{x}^{r}_{j}=\arg\min_{\mathbf{x}\in C_{j}}\sum_{\mathbf{y}\in C_{j}}\|\mathbf{x}-\mathbf{y}\|_{2}$
5:  $D^{r}\leftarrow D^{r}\cup\{x_{i}\in D:\mathbf{x}_{i}=\mathbf{x}^{r}_{j}\}$
6:end for
7: Initialize $\mathbf{Z}^{r}\leftarrow I_{d}$ {one-hot vectors for $D^{r}$}
8:return Initial Representative Texts $D^{r}$, $\mathbf{Z}^{r}$
Algorithm 2 Construct Bag-of-Texts Vectors
[…]
0: Bag-of-texts vectors $\mathbf{Z}\in\mathbb{R}^{N\times(d+d^{*})}$
1: Initialize $\mathbf{z}_{j}\leftarrow\mathbf{0}$ for all $x_{j}\in D\setminus D^{r}$
2:for each $x_{j}\in D\setminus D^{r}$ do
3:  $M_{j}\leftarrow\{x_{j1},\dots,x_{jm}\}\in D^{r}$ from $\mathbf{X}$ as the top-$m$ by cosine similarity to $x_{j}$
4:  $O_{j}\leftarrow\text{LLM}(M_{j})\in\mathcal{P}(M_{j})\cup\{\emptyset\}$ {# $\mathcal{P}(.)$: power set}
5:  if $O_{j}\neq\emptyset$ then
6:   $\mathbf{z}_{j}\leftarrow\text{MeanPooling}\{\mathbf{z}_{k}:x_{k}\in O_{j}\}$
7:   $\mathbf{z}_{j}\leftarrow\frac{\mathbf{z}_{j}}{\|\mathbf{z}_{j}\|_{2}}$
8:  end if
9:end for
[…]
12: Let $\mathbf{P}\leftarrow\{\mathbf{z}_{j}:\mathbf{z}_{j}=\mathbf{0}\}$; one-hot encode each $\mathbf{z}_{j}\in\mathbf{P}$ in new dimensions, and pad the first $d$ columns with zeros.
13: Let $\mathbf{Q}\leftarrow\{\mathbf{z}_{j}:\mathbf{z}_{j}\neq\mathbf{0}\}$; pad $\mathbf{Q}$ and $\mathbf{Z}^{r}$ with $d^{*}$ zero columns
14:return $\mathbf{Z}\leftarrow\text{Stack}(\mathbf{Z}^{r},\mathbf{P},\mathbf{Q})$, rows ordered according to $D$
We first encode all texts $D$ with a pre-trained embedder to obtain embeddings $\mathbf{X}$. To select initial representative texts $D^{r}$, we perform agglomerative clustering (Murtagh and Legendre, 2014) to partition them into $d$ clusters, $d$ being a trivially large number (see Section 5.3). The medoid of each cluster is chosen as a representative and converted into a one-hot vector with initially equal spacing. See Algorithm 1 for specification.
The remaining texts are positioned based on LLM preference. Specifically, for each text $x_{j}\in D\setminus D^{r}$, we retrieve the top $m$ representative texts from the pretrained embeddings $\mathbf{X}$ that are closest based on cosine similarity. An LLM is then used to select which representative(s) are similar. The bag-of-texts vector for each text is computed by averaging the vectors of its selected representatives. After processing all texts, those with no selected representatives are promoted to new representatives and added to $D^{r}$ (so that $|D^{r}|=d+d^{*}$, where $d^{*}$ is the number of such new representatives), and the dimensionality of the BoT vectors is expanded to match. See Algorithm 2 for specification.
This procedure ensures that the initial BoT vectors are evenly spaced for the representative texts, while the vectors for the remaining texts encode LLM-informed relationships, avoiding the need to use the full dimensionality of all texts.
4.2 Iterative stage: bag-of-texts vectors refinements
After constructing the initial positions for all texts, we iteratively update the BoT vectors to better align their relative distances with LLM preferences. The update procedure follows the same approach as in Algorithm 2, with three key modifications. First, the update uses all texts as the pool for subset selection, since each text has a BoT vector. Second, for each text, we select a subset of $m$ candidates based on the concatenation of its pre-trained embedding and BoT vector, $\mathbf{U}^{t}\triangleq\mathbf{X}_{j}\oplus\mathbf{Z}_{j}^{t}$. This ensures that the pre-trained embedding initially guides the BoT vectors, which carry minimal information at the start. Third, the update is performed iteratively rather than just once, allowing the candidate set to vary dynamically as the BoT vectors are refined. The update for a text stops when its BoT vector changes negligibly, as measured by a cosine similarity greater than 0.99. We set $m=30$ and cap the maximum number of iterations at $T=10$ as budget. See Algorithm 3 in Appendix B for specification. Note that the pretrained embeddings $\mathbf{X}$ are used only for updating the Bot vectors; only the final BoT vectors $\mathbf{Z}^{T}$ will be used for clustering.